Sara is 8 years old and her mom is 34 years old. In how many years will Sara be twice as young as her mom will be?

Answer by mathsolverplus(88)

Let x be the number of years,
Sara's age in x years is represented on the left
while her mom's age in x years is represented on the right
Notice how I have added x in order to determine their age in x years respectively:
2(8+x) = 34+x
16+2x = 34+x
-18 = -x
x = 18
Therefore, Sara will be twice as young as her mom in 18 years!

For more additional instant FREE math help, please check out and follow us on social media!!
Instagram (@mathsolver.plus) : https://www.instagram.com/mathsolver.plus/
Twitter (@MathSolverPlus): https://twitter.com/MathSolverPlus
Snapchat (@mathsolverplus)
Email: mathsolverplus@gmail.com